# TaP: A Taxonomy-Guided Framework for Automated and Scalable Preference Data Generation

### v3

@@ Appendix E Experimental Results
Alpaca-GPT-4-ZH (DBLP:journals/corr/abs-2304-03277) 5.07 3.99 4.53 5.35 4.60 5.31 5.11 5.11 5.35 5.14 4.83
COIG (DBLP:journals/corr/abs-2304-07987) 3.35 2.87 3.11 4.57 4.24 4.13 3.87 4.79 4.70 4.38 3.75
MOSS-SFT (MIR-2023-12-294) 5.09 4.01 4.55 5.30 4.47 4.85 5.16 5.76 5.50 5.17 4.86
COIG-CQIA (DBLP:journals/corr/abs-2403-18058) 5.12 3.90 4.51 4.98 5.09 4.81 4.96 5.26 5.01 5.02 4.76
@@
Alpaca-GPT-4-ZH (DBLP:journals/corr/abs-2304-03277) 2.35 2.93 2.64 3.95 3.50 3.88 4.49 4.34 4.56 4.12 3.38
COIG (DBLP:journals/corr/abs-2304-07987) 1.68 2.08 1.88 3.01 2.38 2.78 2.84 3.71 3.52 3.04 2.46
MOSS-SFT (MIR-2023-12-294) 2.51 2.87 2.69 4.10 3.09 4.04 4.57 5.37 4.74 4.32 3.50
COIG-CQIA (DBLP:journals/corr/abs-2403-18058) 2.24 2.53 2.39 3.40 2.93 3.43 3.45 4.18 3.97 3.56 2.97
@@
Alpaca-GPT-4-ZH (DBLP:journals/corr/abs-2304-03277) 3.18 3.21 3.19 4.24 2.76 4.25 4.47 4.74 4.59 4.17 3.68
COIG (DBLP:journals/corr/abs-2304-07987) 1.64 2.26 1.95 2.90 1.95 2.46 2.56 3.42 3.53 2.80 2.38
MOSS-SFT (MIR-2023-12-294) 2.62 2.85 2.73 3.94 3.02 3.70 4.53 5.50 4.46 4.19 3.46
COIG-CQIA (DBLP:journals/corr/abs-2403-18058) 2.49 2.79 2.64 3.26 2.74 3.75 3.45 4.00 4.14 3.56 3.10
@@
Alpaca-GPT-4-ZH (DBLP:journals/corr/abs-2304-03277) 5.89 4.79 5.34 5.77 5.41 5.88 5.36 6.05 5.90 5.73 5.54
COIG (DBLP:journals/corr/abs-2304-07987) 3.50 3.25 3.38 4.91 4.62 4.95 4.41 5.61 4.68 4.86 4.12
MOSS-SFT (MIR-2023-12-294) 5.03 4.89 4.96 5.47 4.79 5.15 5.71 5.89 5.89 5.48 5.22
COIG-CQIA (DBLP:journals/corr/abs-2403-18058) 5.40 4.47 4.94 5.48 5.00 5.55 5.17 5.13 5.21 5.26 5.10
@@
Qwen2.5-3B-SFT-Open Huozi-RLHF (huozi) 4.63 3.42 4.02 4.86 4.74 4.87 5.55 5.61 5.74 5.23 4.63
@@
Qwen2.5-3B-SFT-TaP Huozi-RLHF (huozi) 6.01 4.29 5.15 5.64 5.49 5.37 6.23 6.21 5.98 5.82 5.49
@@
Qwen2.5-7B-SFT-Open Huozi-RLHF (huozi) 5.04 4.01 4.53 5.48 5.02 5.24 5.43 5.68 5.50 5.39 4.96
@@
Qwen2.5-7B-SFT-TaP Huozi-RLHF (huozi) 7.44 5.04 6.24 6.31 5.78 6.34 6.41 6.79 6.39 6.34 6.29
@@
Llama-3.1-8B-SFT-Open Huozi-RLHF (huozi) 3.68 3.37 3.52 4.98 4.43 4.78 5.40 5.74 5.83 5.19 4.36
@@
Llama-3.1-8B-SFT-TaP Huozi-RLHF (huozi) 3.47 3.93 3.70 4.84 3.93 4.87 5.71 6.18 5.78 5.22 4.46
@@
Gemma-2-9B-SFT-Open Huozi-RLHF (huozi) 4.36 3.51 3.93 5.01 4.10 5.15 5.57 5.55 5.93 5.22 4.58
@@
Gemma-2-9B-SFT-TaP Huozi-RLHF (huozi) 4.78 3.59 4.18 4.94 3.83 5.09 5.71 5.95 5.84 5.23 4.70
@@
Qwen2.5-14B-SFT-Open Huozi-RLHF (huozi) 6.17 5.34 5.75 6.19 5.74 6.18 5.61 5.76 5.90 5.90 5.82
@@
Qwen2.5-14B-SFT-TaP Huozi-RLHF (huozi) 7.54 6.35 6.94 6.95 6.38 6.75 6.55 6.68 6.51 6.64 6.79
@@
Qwen2.5-3B-SFT-Open Huozi-RLHF (huozi) 4.45 3.30 3.88 4.75 4.21 4.72 5.48 5.45 5.64 5.04 4.46
@@
Qwen2.5-3B-SFT-TaP Huozi-RLHF (huozi) 5.45 3.85 4.65 5.45 5.09 5.33 5.87 5.55 5.74 5.50 5.08
@@
Qwen2.5-7B-SFT-Open Huozi-RLHF (huozi) 4.86 3.88 4.37 5.52 4.66 5.19 5.43 5.84 5.40 5.34 4.85
@@
Qwen2.5-7B-SFT-TaP Huozi-RLHF (huozi) 6.79 4.88 5.83 5.86 5.36 6.34 6.09 6.11 6.26 6.00 5.92
@@
Llama-3.1-8B-SFT-Open Huozi-RLHF (huozi) 3.29 3.29 3.29 4.71 3.97 4.65 5.52 5.63 5.88 5.06 4.17
@@
Llama-3.1-8B-SFT-TaP Huozi-RLHF (huozi) 3.01 3.64 3.33 4.61 3.22 4.65 5.39 5.61 5.71 4.86 4.09
@@
Gemma-2-9B-SFT-Open Huozi-RLHF (huozi) 3.72 3.35 3.54 4.81 3.72 4.99 5.49 5.39 5.90 5.05 4.29
@@
Gemma-2-9B-SFT-TaP Huozi-RLHF (huozi) 4.37 3.17 3.77 4.69 3.43 5.25 5.32 5.34 5.56 4.93 4.35
@@
Qwen2.5-14B-SFT-Open Huozi-RLHF (huozi) 6.12 5.17 5.64 5.98 5.64 5.90 5.51 5.87 5.97 5.81 5.73
@@
Qwen2.5-14B-SFT-TaP Huozi-RLHF (huozi) 7.19 5.64 6.41 6.41 5.83 6.60 6.19 6.24 6.43 6.28 6.35
@@
Qwen2.5-3B-SFT-Open Huozi-RLHF (huozi) 4.62 3.65 4.13 4.45 4.71 4.57 5.03 5.18 5.36 4.88 4.51
@@
Qwen2.5-3B-SFT-TaP Huozi-RLHF (huozi) 5.59 3.63 4.61 4.97 4.79 5.07 5.48 5.79 5.53 5.27 4.94
@@
Qwen2.5-7B-SFT-Open Huozi-RLHF (huozi) 5.04 4.39 4.71 5.60 5.00 5.68 5.61 5.92 5.53 5.56 5.13
@@
Qwen2.5-7B-SFT-TaP Huozi-RLHF (huozi) 7.11 5.38 6.24 6.45 6.19 6.26 6.40 6.89 6.46 6.44 6.34
@@
Llama-3.1-8B-SFT-Open Huozi-RLHF (huozi) 3.78 3.66 3.72 5.02 4.36 4.91 5.67 5.92 5.74 5.27 4.50
@@
Llama-3.1-8B-SFT-TaP Huozi-RLHF (huozi) 3.81 4.08 3.94 4.81 4.21 5.09 5.73 5.89 5.78 5.25 4.60
@@
Gemma-2-9B-SFT-Open Huozi-RLHF (huozi) 4.23 3.26 3.75 5.09 3.91 4.44 4.88 5.03 5.16 4.75 4.25
@@
Gemma-2-9B-SFT-TaP Huozi-RLHF (huozi) 4.13 3.57 3.85 5.04 4.53 5.13 5.53 5.26 5.73 5.20 4.53
@@
Qwen2.5-14B-SFT-Open Huozi-RLHF (huozi) 6.03 5.18 5.61 5.50 6.05 6.35 5.73 5.82 5.93 5.90 5.75
@@
Qwen2.5-14B-SFT-TaP Huozi-RLHF (huozi) 6.79 6.43 6.61 6.07 6.38 6.22 6.59 7.03 6.69 6.50 6.55
@@
Qwen2.5-3B-SFT-Open Huozi-RLHF (huozi) 4.93 3.96 4.44 5.10 4.81 4.72 5.68 5.76 6.07 5.36 4.90
@@
Qwen2.5-3B-SFT-TaP Huozi-RLHF (huozi) 5.33 3.90 4.62 5.28 4.95 5.37 5.81 5.89 6.00 5.55 5.08
@@
Qwen2.5-7B-SFT-Open Huozi-RLHF (huozi) 4.85 4.08 4.46 5.19 4.55 5.60 5.67 5.84 5.72 5.43 4.95
@@
Qwen2.5-7B-SFT-TaP Huozi-RLHF (huozi) 6.67 4.67 5.67 6.04 5.52 5.69 6.03 6.26 6.32 5.98 5.82
@@
Llama-3.1-8B-SFT-Open Huozi-RLHF (huozi) 3.54 3.27 3.41 4.81 3.79 4.43 5.16 5.82 5.80 4.97 4.19
@@
Llama-3.1-8B-SFT-TaP Huozi-RLHF (huozi) 3.50 3.73 3.61 4.28 3.66 4.53 5.36 5.50 5.81 4.86 4.24
@@
Gemma-2-9B-SFT-Open Huozi-RLHF (huozi) 3.98 3.63 3.81 4.98 4.00 4.99 5.25 5.92 5.71 5.14 4.47
@@
Gemma-2-9B-SFT-TaP Huozi-RLHF (huozi) 3.62 3.50 3.56 4.69 3.83 4.96 5.17 5.50 5.61 4.96 4.26
@@
Qwen2.5-14B-SFT-Open Huozi-RLHF (huozi) 6.53 4.86 5.69 5.83 5.71 6.25 5.64 6.03 6.10 5.93 5.81
@@
Qwen2.5-14B-SFT-TaP Huozi-RLHF (huozi) 6.94 5.90 6.42 6.55 5.88 6.58 6.17 6.34 6.49 6.34 6.38

### v4

@@ Appendix E Experimental Results
Alpaca-GPT-4-ZH (Peng et al., 2023) 5.07 3.99 4.53 5.35 4.60 5.31 5.11 5.11 5.35 5.14 4.83
COIG (Zhang et al., 2023a) 3.35 2.87 3.11 4.57 4.24 4.13 3.87 4.79 4.70 4.38 3.75
MOSS-SFT (Sun et al., 2024) 5.09 4.01 4.55 5.30 4.47 4.85 5.16 5.76 5.50 5.17 4.86
COIG-CQIA (Bai et al., 2024) 5.12 3.90 4.51 4.98 5.09 4.81 4.96 5.26 5.01 5.02 4.76
@@
Alpaca-GPT-4-ZH (Peng et al., 2023) 2.35 2.93 2.64 3.95 3.50 3.88 4.49 4.34 4.56 4.12 3.38
COIG (Zhang et al., 2023a) 1.68 2.08 1.88 3.01 2.38 2.78 2.84 3.71 3.52 3.04 2.46
MOSS-SFT (Sun et al., 2024) 2.51 2.87 2.69 4.10 3.09 4.04 4.57 5.37 4.74 4.32 3.50
COIG-CQIA (Bai et al., 2024) 2.24 2.53 2.39 3.40 2.93 3.43 3.45 4.18 3.97 3.56 2.97
@@
Alpaca-GPT-4-ZH (Peng et al., 2023) 3.18 3.21 3.19 4.24 2.76 4.25 4.47 4.74 4.59 4.17 3.68
COIG (Zhang et al., 2023a) 1.64 2.26 1.95 2.90 1.95 2.46 2.56 3.42 3.53 2.80 2.38
MOSS-SFT (Sun et al., 2024) 2.62 2.85 2.73 3.94 3.02 3.70 4.53 5.50 4.46 4.19 3.46
COIG-CQIA (Bai et al., 2024) 2.49 2.79 2.64 3.26 2.74 3.75 3.45 4.00 4.14 3.56 3.10
@@
Alpaca-GPT-4-ZH (Peng et al., 2023) 5.89 4.79 5.34 5.77 5.41 5.88 5.36 6.05 5.90 5.73 5.54
COIG (Zhang et al., 2023a) 3.50 3.25 3.38 4.91 4.62 4.95 4.41 5.61 4.68 4.86 4.12
MOSS-SFT (Sun et al., 2024) 5.03 4.89 4.96 5.47 4.79 5.15 5.71 5.89 5.89 5.48 5.22
COIG-CQIA (Bai et al., 2024) 5.40 4.47 4.94 5.48 5.00 5.55 5.17 5.13 5.21 5.26 5.10
@@
Qwen2.5-3B-SFT-Open Huozi-RLHF (Huozi-Team, 2024) 4.63 3.42 4.02 4.86 4.74 4.87 5.55 5.61 5.74 5.23 4.63
@@
Qwen2.5-3B-SFT-TaP Huozi-RLHF (Huozi-Team, 2024) 6.01 4.29 5.15 5.64 5.49 5.37 6.23 6.21 5.98 5.82 5.49
@@
Qwen2.5-7B-SFT-Open Huozi-RLHF (Huozi-Team, 2024) 5.04 4.01 4.53 5.48 5.02 5.24 5.43 5.68 5.50 5.39 4.96
@@
Qwen2.5-7B-SFT-TaP Huozi-RLHF (Huozi-Team, 2024) 7.44 5.04 6.24 6.31 5.78 6.34 6.41 6.79 6.39 6.34 6.29
@@
Llama-3.1-8B-SFT-Open Huozi-RLHF (Huozi-Team, 2024) 3.68 3.37 3.52 4.98 4.43 4.78 5.40 5.74 5.83 5.19 4.36
@@
Llama-3.1-8B-SFT-TaP Huozi-RLHF (Huozi-Team, 2024) 3.47 3.93 3.70 4.84 3.93 4.87 5.71 6.18 5.78 5.22 4.46
@@
Gemma-2-9B-SFT-Open Huozi-RLHF (Huozi-Team, 2024) 4.36 3.51 3.93 5.01 4.10 5.15 5.57 5.55 5.93 5.22 4.58
@@
Gemma-2-9B-SFT-TaP Huozi-RLHF (Huozi-Team, 2024) 4.78 3.59 4.18 4.94 3.83 5.09 5.71 5.95 5.84 5.23 4.70
@@
Qwen2.5-14B-SFT-Open Huozi-RLHF (Huozi-Team, 2024) 6.17 5.34 5.75 6.19 5.74 6.18 5.61 5.76 5.90 5.90 5.82
@@
Qwen2.5-14B-SFT-TaP Huozi-RLHF (Huozi-Team, 2024) 7.54 6.35 6.94 6.95 6.38 6.75 6.55 6.68 6.51 6.64 6.79
@@
Qwen2.5-3B-SFT-Open Huozi-RLHF (Huozi-Team, 2024) 4.45 3.30 3.88 4.75 4.21 4.72 5.48 5.45 5.64 5.04 4.46
@@
Qwen2.5-3B-SFT-TaP Huozi-RLHF (Huozi-Team, 2024) 5.45 3.85 4.65 5.45 5.09 5.33 5.87 5.55 5.74 5.50 5.08
@@
Qwen2.5-7B-SFT-Open Huozi-RLHF (Huozi-Team, 2024) 4.86 3.88 4.37 5.52 4.66 5.19 5.43 5.84 5.40 5.34 4.85
@@
Qwen2.5-7B-SFT-TaP Huozi-RLHF (Huozi-Team, 2024) 6.79 4.88 5.83 5.86 5.36 6.34 6.09 6.11 6.26 6.00 5.92
@@
Llama-3.1-8B-SFT-Open Huozi-RLHF (Huozi-Team, 2024) 3.29 3.29 3.29 4.71 3.97 4.65 5.52 5.63 5.88 5.06 4.17
@@
Llama-3.1-8B-SFT-TaP Huozi-RLHF (Huozi-Team, 2024) 3.01 3.64 3.33 4.61 3.22 4.65 5.39 5.61 5.71 4.86 4.09
@@
Gemma-2-9B-SFT-Open Huozi-RLHF (Huozi-Team, 2024) 3.72 3.35 3.54 4.81 3.72 4.99 5.49 5.39 5.90 5.05 4.29
@@
Gemma-2-9B-SFT-TaP Huozi-RLHF (Huozi-Team, 2024) 4.37 3.17 3.77 4.69 3.43 5.25 5.32 5.34 5.56 4.93 4.35
@@
Qwen2.5-14B-SFT-Open Huozi-RLHF (Huozi-Team, 2024) 6.12 5.17 5.64 5.98 5.64 5.90 5.51 5.87 5.97 5.81 5.73
@@
Qwen2.5-14B-SFT-TaP Huozi-RLHF (Huozi-Team, 2024) 7.19 5.64 6.41 6.41 5.83 6.60 6.19 6.24 6.43 6.28 6.35
@@
Qwen2.5-3B-SFT-Open Huozi-RLHF (Huozi-Team, 2024) 4.62 3.65 4.13 4.45 4.71 4.57 5.03 5.18 5.36 4.88 4.51
@@
Qwen2.5-3B-SFT-TaP Huozi-RLHF (Huozi-Team, 2024) 5.59 3.63 4.61 4.97 4.79 5.07 5.48 5.79 5.53 5.27 4.94
@@
Qwen2.5-7B-SFT-Open Huozi-RLHF (Huozi-Team, 2024) 5.04 4.39 4.71 5.60 5.00 5.68 5.61 5.92 5.53 5.56 5.13
@@
Qwen2.5-7B-SFT-TaP Huozi-RLHF (Huozi-Team, 2024) 7.11 5.38 6.24 6.45 6.19 6.26 6.40 6.89 6.46 6.44 6.34
@@
Llama-3.1-8B-SFT-Open Huozi-RLHF (Huozi-Team, 2024) 3.78 3.66 3.72 5.02 4.36 4.91 5.67 5.92 5.74 5.27 4.50
@@
Llama-3.1-8B-SFT-TaP Huozi-RLHF (Huozi-Team, 2024) 3.81 4.08 3.94 4.81 4.21 5.09 5.73 5.89 5.78 5.25 4.60
@@
Gemma-2-9B-SFT-Open Huozi-RLHF (Huozi-Team, 2024) 4.23 3.26 3.75 5.09 3.91 4.44 4.88 5.03 5.16 4.75 4.25
@@
Gemma-2-9B-SFT-TaP Huozi-RLHF (Huozi-Team, 2024) 4.13 3.57 3.85 5.04 4.53 5.13 5.53 5.26 5.73 5.20 4.53
@@
Qwen2.5-14B-SFT-Open Huozi-RLHF (Huozi-Team, 2024) 6.03 5.18 5.61 5.50 6.05 6.35 5.73 5.82 5.93 5.90 5.75
@@
Qwen2.5-14B-SFT-TaP Huozi-RLHF (Huozi-Team, 2024) 6.79 6.43 6.61 6.07 6.38 6.22 6.59 7.03 6.69 6.50 6.55
@@
Qwen2.5-3B-SFT-Open Huozi-RLHF (Huozi-Team, 2024) 4.93 3.96 4.44 5.10 4.81 4.72 5.68 5.76 6.07 5.36 4.90
@@
Qwen2.5-3B-SFT-TaP Huozi-RLHF (Huozi-Team, 2024) 5.33 3.90 4.62 5.28 4.95 5.37 5.81 5.89 6.00 5.55 5.08
@@
Qwen2.5-7B-SFT-Open Huozi-RLHF (Huozi-Team, 2024) 4.85 4.08 4.46 5.19 4.55 5.60 5.67 5.84 5.72 5.43 4.95
@@
Qwen2.5-7B-SFT-TaP Huozi-RLHF (Huozi-Team, 2024) 6.67 4.67 5.67 6.04 5.52 5.69 6.03 6.26 6.32 5.98 5.82
@@
Llama-3.1-8B-SFT-Open Huozi-RLHF (Huozi-Team, 2024) 3.54 3.27 3.41 4.81 3.79 4.43 5.16 5.82 5.80 4.97 4.19
@@
Llama-3.1-8B-SFT-TaP Huozi-RLHF (Huozi-Team, 2024) 3.50 3.73 3.61 4.28 3.66 4.53 5.36 5.50 5.81 4.86 4.24
@@
Gemma-2-9B-SFT-Open Huozi-RLHF (Huozi-Team, 2024) 3.98 3.63 3.81 4.98 4.00 4.99 5.25 5.92 5.71 5.14 4.47
@@
Gemma-2-9B-SFT-TaP Huozi-RLHF (Huozi-Team, 2024) 3.62 3.50 3.56 4.69 3.83 4.96 5.17 5.50 5.61 4.96 4.26
@@
Qwen2.5-14B-SFT-Open Huozi-RLHF (Huozi-Team, 2024) 6.53 4.86 5.69 5.83 5.71 6.25 5.64 6.03 6.10 5.93 5.81
@@
Qwen2.5-14B-SFT-TaP Huozi-RLHF (Huozi-Team, 2024) 6.94 5.90 6.42 6.55 5.88 6.58 6.17 6.34 6.49 6.34 6.38